OpenStudy (anonymous):

how can i simplify this (x^2 - 8x + 15) / (x - 5)

OpenStudy (anonymous):

Can you factor the numerator?

OpenStudy (anonymous):

i dnt know how to factor the numerator

OpenStudy (anonymous):

Well, what two numbers add up to -8 but multiply together to get 15?

OpenStudy (anonymous):

no teo numbers multiply together to get 15

OpenStudy (anonymous):

hint one of them is -3 -3 + x = -8 -3 * x = 15

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

yeah, so you factor the numerator as (x-5)(x-3), now can you simplify the expression?

OpenStudy (anonymous):

i dont get how to simplify

OpenStudy (anonymous):

Ok, its kind of like .. (5*3) / 5, how do you simplify that?

OpenStudy (anonymous):

the answer would be 3

OpenStudy (anonymous):

so the answer would be x-3

OpenStudy (anonymous):

Yeah, because you can just cross out the 5's, because you multiply then divide by 5 Its the same thing here: (x-5)(x-3) ------- (x-5) Any idea how to do it now?

OpenStudy (anonymous):

so the answer would be x-3 ??

OpenStudy (anonymous):

Yup :), get it?

OpenStudy (anonymous):

ya thanks so much